Find these terms of the sequence {an}, where an =




2 · (−3)n + 5n.




a) a0




b) a1




c) a4




d) a5






1
Expert's answer
2021-12-13T14:43:42-0500

Let us find these terms of the sequence "(a_n)," where "a_n = 2 \u00b7 (\u22123)^n + 5n."


a) It follows that "a_0 = 2 · (−3)^0 + 5\cdot 0=2."


b) It follows that "a_1 = 2 · (−3)^1 + 5\cdot 1=-6+5=-1."


c) It follows that "a_4 = 2 · (−3)^4 + 5\cdot 4=162+20=182."


d) It follows that "a_5 = 2 · (−3)^5 + 5\cdot 5=-486+25=-461."